OpenStudy (welshfella):

I guess i've done problems like this in the past but I'm struggling with this one:- The polynomial Q(x) leaves remainder 4 when divided by x - 1, and remainder 8 when divided by x + 1. The remainder when Q(x) is divided by x^2 - 1 is A 32 B -4x + 9 C -4x - 7

OpenStudy (welshfella):

obviously Q(1) = 4 and Q(-1) = 8 by the remainder theorem

OpenStudy (sohailiftikhar):

find the two values of x

OpenStudy (sohailiftikhar):

x^2-1=(x-1)(x+1) so x=1 and x=-1 put one by one in equation

OpenStudy (sohailiftikhar):

so the reminder is 32

OpenStudy (welshfella):

I dont follow that...

OpenStudy (sohailiftikhar):

as given in data reminder for (x-1)=4 and for (x+1)=8 so for x^2-1=(x-1)(x+1)=4*8=32

OpenStudy (sohailiftikhar):

now get it ?

OpenStudy (welshfella):

I don't think that's correct.

OpenStudy (welshfella):

I'd dont think this is particularly different - I'm just missing something

geerky42 (geerky42):

*

OpenStudy (sohailiftikhar):

O.o so what you think huh ? It is correct bro

OpenStudy (sohailiftikhar):

so you just confused get calm and think on it for a minute

OpenStudy (welshfella):

I dont' know - but I don't think your logic is correct

OpenStudy (welshfella):

I'm going to look up the answer . I am confused - you are right there!! lol

OpenStudy (welshfella):

its -4x + 9

OpenStudy (sohailiftikhar):

yea! go an look the answer perhaps then you will believe on my answer

OpenStudy (sohailiftikhar):

no way it can't be

OpenStudy (sohailiftikhar):

take a screen short

OpenStudy (welshfella):

I'm helping my grandson with his maths revision. Well that's the answer in the book.

OpenStudy (welshfella):

I haven't got a scanner

OpenStudy (sohailiftikhar):

it's very simple .. ok tell me how they get 4 when they divide equation by (x-1) huh ?

OpenStudy (sohailiftikhar):

ok ganesh is here he can justify better now

OpenStudy (welshfella):

4 is the remainder and = q(1)

ganeshie8 (ganeshie8):

Firstly, notice that we get a polynomial as remainder that is one degree less than whatever we're dividing by

Nnesha (nnesha):

^

OpenStudy (welshfella):

right

ganeshie8 (ganeshie8):

for example, (x^5+2x+1)/(x^2-1) gives a remainder that looks like \(ax+b\) yes ?

OpenStudy (welshfella):

right

ganeshie8 (ganeshie8):

similarly (x^100 + x+1)/(x^10 + 1) gives a remainder that looks like \(ax^9+bx^8+\cdots\)

ganeshie8 (ganeshie8):

the degree of remainder is always one less than the degree of bottom

OpenStudy (welshfella):

so the remainder in this case must be of the form ax + b?

ganeshie8 (ganeshie8):

right, so lets suppose \[Q(x) = F(x)*(x^2-1)+\color{red}{ax+b}\] our goal is to find that red part

OpenStudy (welshfella):

ok

ganeshie8 (ganeshie8):

since we know that \(Q(1)=4\) and \(Q(-1)=8\), plug them in and get two equations

OpenStudy (sohailiftikhar):

ganesh what you said about the given that reminder of that equation is 4 when divided by (x-1) where is x term with 4 ?

ganeshie8 (ganeshie8):

\[Q(1) = F(1)*(1^2-1)+\color{red}{a*1+b} \implies 4 = \color{red}{a+b} \tag{1}\] \[Q(-1) = F(-1)*((-1)^2-1)+\color{red}{a(-1)+b} \implies 8 = \color{red}{-a+b} \tag{2}\] two equations and two unknowns, we can solve them

ganeshie8 (ganeshie8):

for that, we may think that the coefficient of x is 0 @sohailiftikhar

OpenStudy (welshfella):

thats really clever Thanx ganesh

ganeshie8 (ganeshie8):

np, im getting the remainder is \(-2x+6\) looks the options are wrong

OpenStudy (welshfella):

yes i got that too b = 6 ans a = -2

OpenStudy (welshfella):

I'll just recheck the answer in the book

OpenStudy (sohailiftikhar):

lol

OpenStudy (welshfella):

Yes thats the answer in the book. Well mistakes are made

ganeshie8 (ganeshie8):

happens... our method is pretty robust and straightforward, nothing that could go wrong..

OpenStudy (sohailiftikhar):

from which grades book u got that problem bro ?

OpenStudy (welshfella):

Oh its a pretty old UK Advanced Level book from 1979. Examinations have become a little easier since then. Its wriiten by a professor of Mathematics but mistakes are made by everyone...

OpenStudy (sohailiftikhar):

lol ok

OpenStudy (welshfella):

Open study is a great place to study . There is a wealth a talent here.

OpenStudy (welshfella):

* wealth of talent

OpenStudy (sohailiftikhar):

yes:)